# Simultaneous false discovery bounds for invariant causal prediction

### v2

@@ 3.2 Simultaneous false discovery bounds for ICP
and p-value
@@
By using a similar argument as in Proposition 3.1, one can see that $p^{*}_{S}$ is valid.
@@
For any set $R\subseteq[m]$, let $t_{\alpha}(R)$ be the size of the largest subset of $R$ whose corresponding p-value is larger than $\alpha$. That is,
$t_{\alpha}(R)=\max\{|I|:I\subseteq R\text{ and }p^{*}_{I}>\alpha\}.$ (12)
Then, $t_{\alpha}(R)$ is a simultaneous false discovery upper bound, as shown in the following Theorem.
@@
$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$ (13)
@@
Based on the simultaneous guarantee (13), one can freely check any set $R$ of interest, and a high probability upper bound (12) on the false discoveries can be calculated. Equivalently, $|R|-t_{\alpha}(R)$ is a simultaneous true discovery lower bound:
$P(|R\cap S^{*}|\geq|R|-t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$ (14)
The simultaneous upper bound (12) is the same as the upper bound proposed by Goeman and Solari, (2011) for closed testing (Marcus et al.,, 1976). In fact, our method can be seen as a special case of closed testing. Compared to the standard closed testing procedure, the difference is that no closed testing adjustments are needed in our case. This is due to the specific form of $p^{*}_{S}$. In particular, if $p^{*}_{S}\leq\alpha$, we must have $p^{*}_{S^{\prime}}\leq\alpha$ for all $S\subseteq S^{\prime}$. Thus, if a hypothesis is locally rejected, it must be rejected by closed testing, so there is no need to implement closed testing adjustments.
At last, we mention that calculating the false discovery bound (12) can be computationally intractable for large $|R|$ (or large $m$). In fact, computation is also a fundamental issue for the original ICP approach, which requires testing $2^{|m|}$ hypotheses. Some tricks are suggested to deal with this issue, including first reducing the size of the potential causal predictors by using a pre-screening method such as Lasso. See Peters et al., (2016) for more discussions.
3.3 A discussion about false discovery rate control for ICP
With a multiple testing formulation (6) and valid p-values (7), an alternative strategy to obtain a less conservative result than $\widehat{S}^{\text{ICP}}(\mathcal{E})$ is to control the commonly used false discovery rate (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001; Lehmann and Romano,, 2022). Due to Proposition 3.3, however, directly comparing $p^{*}_{1},\dots,p^{*}_{m}$ to the significance level already controls the familywise error rate, so there is no point in further applying false discovery rate control procedures on these p-values. This observation is interesting because, typically, multiple testing corrections (e.g., Bonferroni correction) on raw p-values are required to control the familywise error rate. Thus, it is natural to ask: Can we obtain smaller valid p-values than $p^{*}_{1},\dots,p^{*}_{m}$? If so, one may apply false discovery rate control procedures to obtain more discoveries than ICP. However, in the following, we argue that this approach does not seem promising.
Specifically, we show that if some $p^{\prime}_{i}$ satisfying
${\mathbb{P}}(p^{\prime}_{i}\leq p^{*}_{i})=1\quad\text{and}\quad{\mathbb{P}}_{H^{*}_{0,i}}(p^{\prime}_{i}<p^{*}_{i})>0,$ (15)
then it is not valid in general. In particular, suppose that we have an ideal test for $H_{0,S}(\mathcal{E})$ that yields p-values $p_{S}\sim U(0,1)$ for true nulls and $p_{S}=0$ for false nulls. Consider a setting where the only true null hypothesis is $H_{0,S^{*}}(\mathcal{E})$. This is the desired setting where all causal predictors are identifiable, and it can happen, for example, when the environments are rich enough. The following proposition shows that $p^{\prime}_{i}$ satisfying (15) is not valid in such settings.
Proposition 3.5.
Let $S^{*}$ be the set of true causal predictors. Consider the desired case where only $H_{0,S^{*}}(\mathcal{E})$ is true, $p_{S^{*}}\sim U(0,1)$ and $p_{S}=0$ for any $S\neq S^{*}$. Then, for $p^{\prime}_{i}$ satisfying (15), there exists some $c^{*}\in(0,1)$ such that $P_{H^{*}_{0,i}}(p^{\prime}_{i}\leq c^{*})>c^{*}$.
This result excludes the possibility of constructing $p_{i}^{\prime}$ based on some analytic function, such that $p_{i}^{\prime}$ is strictly smaller than $p_{i}^{*}$ in probability $1$ (recall that $p_{i}^{*}$ is obtained by a maximum function (7)). Consequently, considering false discovery rate control for ICP does not appear to be a promising idea.
@@ 4.1 Numerical simulations
We implement simulations to verify the simultaneous guarantee (13) empirically and to illustrate the extra information one may obtain by using the simultaneous false discovery bound compared to the original ICP approach. All simulations were carried out in R, and the code is available at https://github.com/Jinzhou-Li/ICPsimultaneousBounds.
@@
$X^{e}\leftarrow BX^{e}+\epsilon^{e},$
where $X^{e}=(X^{e}_{1},\dots,X^{e}_{10})^{T}$, $B$ is a strict lower-triangular matrix, $\epsilon^{e}\sim N_{10}(b^{e},\text{diag}(\sigma))$ and $e\in\{1,2,3,4,5\}$. That is, we consider mean-shift intervention and five environments. We treat $X_{6}$ as the response in our simulations.
We implement $500$ simulations. In each simulation, we randomly sample the non-zero elements of matrix $B$ with a non-zero probability $0.8$. Their values are then uniformly sampled from $1$ to $2$. Each entry of $\sigma$ is uniformly sampled from $0.5$ to $1.5$. For the mean-shift intervention, we set $b^{1}=0$. For other $b_{i}$, we randomly select $3$ non-zero entries excluding $X_{6}$. That is, we consider $3$ intervened variables. The non-zero entries of $b_{i}$ are sampled from $N(10,5^{2})$. We generate $100$ samples for each environment and use a significance level $\alpha=0.05$.
For each generated dataset, we apply ICP and record the number of discoveries. Based on the same p-values used by ICP, we calculate the simultaneous false discovery upper bound $t_{\alpha}(R)$ (see (12)) and the true discovery lower bound $|R|-t_{\alpha}$ (see (14)) for all $5122^{9}=$ 512 sets.
@@
Over these $500$ simulations, the empirical probability of $P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])$ is $0.988$, which is larger than $0.951-\alpha=$ 0.95, so the simultaneous guarantee (13) empirically holds. The average number of causal predictors is $4.144$, and the average number of discoveries of ICP is $0.52$. The first plot in Figure 1 shows the average false discovery upper bounds and true discovery lower bounds for all $512$ sets, as well as the size of each set. Note that the x-axis is only the index for each set, and we didn’t present which sets these indices refer to for simplicity. Compared to ICP, extra information can be obtained by looking at these bounds. For example, by looking at the true discovery lower bound of the set containing all predictors (with set index $512$), we know that, on average, there are about three true causal predictors in all nine predictors.
To better illustrate the usage of the simultaneous bounds in practice, we look at the results of two simulated datasets rather than the averaged results. These results are shown in the second and third plots of Figure 1, respectively.
For the first simulated dataset, the true causal predictors are $X_{1},X_{2},X_{3}$ and $X_{4}$. ICP discovers $X_{2}$ and $X_{3}$, meaning that with a probability larger than $0.95$, $X_{2}$ and $X_{3}$ are causal predictors. That is all the information we can get by using ICP. By looking at the simultaneous bounds (see the second plot in Figure 1), however, extra information can be obtained. For example, we can see that the sets with indices $131$ and $257$ has true discovery lower bounds $3$ and $4$, respectively. These two sets are $\{X_{1},X_{2},X_{3},X_{4}\}$ and $\{X_{1},X_{2},X_{3},X_{4},X_{5}\}$. Hence, we know that with probability larger than $0.95$, there are at least three causal predictors in $\{X_{1},X_{2},X_{3},X_{4}\}$, and at least four causal predictors in $\{X_{1},X_{2},X_{3},X_{4},X_{5}\}$. Note that the same information that both $X_{2}$ and $X_{3}$ are true causal predictors can also be obtained, because the true discovery lower bound for $\{X_{2},X_{3}\}$ (with set index $13$) is two.
@@ 4.2 Real data application
We look at a real dataset used in Stock et al., (2003) about the educational attainment of teenagers (Rouse,, 1995). This dataset contains information of $4739$ students from approximately 1100 US high schools, including gender, ethnicity, achievement test score, whether the father or mother is a college graduate, whether the family owns their home, whether the school is in an urban area, county unemployment rate, state hourly wage in manufacturing, average college tuition, whether the family income above 25000 per year, and region.
@@
We first apply ICP with a significance level $\alpha=0.05$, returning one variable ‘score’ as the causal predictor. It seems reasonable that test score has a causal influence on whether the student would attain a BA degree or higher. We then calculate the simultaneous bounds for all $81922^{13}=$ 8192 sets, and the results are shown in Figure 2. More information can be obtained by looking at these bounds. For example, by looking at the set with index $112$, we know that with a probability larger than $0.95$, there are at least two causal predictors in {‘score’, ’fcollege__\__no’, ’mcollege__\__no’}. It seems plausible that the true discovery lower bound is $2$ for this set, because the two variables of whether father or mother went to college are highly correlated, so it is difficult to distinguish between whether both are causal predictors or only one of them is.
@@ Acknowledgement
The author thanks Jelle Goeman and Nicolai Meinshausen for helpful discussions and suggestions on the draft of this paper. The author gratefully acknowledges support by the SNSF Grant P500PT-210978.
Appendix A Supplementary material
A.1 Proof of Proposition 1
Proposition A.1.
For any $c\in(0,1)$, $P_{H^{*}_{0,i}}(p^{*}_{i}\leq c)\leq c$.
@@ A.1 Proof of Proposition 1
When $H^{*}_{0,i}$ is true, we have $S^{*}\subseteq[m]\setminus\{i\}$ as $i\not\in S^{*}$, so $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}$. Therefore, for any $c\in(0,1)$,
@@
A.2 Proof of Proposition 2
Proposition A.2.
[…]
To prove Proposition 2, we first introduce the following lemma, which gives the relation between $\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}(\alpha)\leq\alpha\}$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$.
Lemma A.1.
[…]
We first prove Lemma A.1.
@@ A.2 Proof of Proposition 2
If $\widehat{S}(\mathcal{E})=\emptyset$, that is, $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}>\alpha$ for all $i\in[m]$, then we must have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$. Otherwise, if there exists some $j\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, then for any $R$ whose corresponding $p_{R}>\alpha$, we must have $j\in R$ by the definition of $\widehat{S}^{\text{ICP}}(\mathcal{E})$. But $p^{*}_{j}=\max_{S\subseteq[m]\setminus\{j\}}p_{S}>\alpha$ implies that there exists some $R^{\prime}$ not containing $j$ and $p_{R^{\prime}}>\alpha$, which leads to a contradiction. If $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$, by the assumption in the proposition, there must exist some $R_{1}$ and $R_{2}$ such that $p_{R_{1}}>\alpha$, $p_{R_{2}}>\alpha$ and $R_{1}\cap R_{2}=\emptyset$. Then we must have $\widehat{S}(\mathcal{E})=\emptyset$. Otherwise, if there exists some $j\in\widehat{S}(\mathcal{E})$, that is, $p^{*}_{j}=\max_{S\subseteq[m]\setminus\{j\}}p_{S}\leq\alpha$, then in order for $p_{R_{1}}>\alpha$ and $p_{R_{2}}>\alpha$, both $R_{1}$ and $R_{2}$ must contain $j$, which contradicts the fact that $R_{1}\cap R_{2}=\emptyset$. Now assume both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ are non-empty. If $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, that is, for any $S$ such that $H_{0,S}(\mathcal{E})$ is not rejected (that is, $p_{S}>\alpha$), we have $i\in S$. So for any $R\subseteq[m]\setminus\{i\}$, $H_{0,R}(\mathcal{E})$ must be rejected, so $p^{*}_{i}=\max_{R\subseteq[m]\setminus\{i\}}p_{R}\leq\alpha$, which implies that $i\in\widehat{S}(\mathcal{E})$. If $i\in\widehat{S}(\mathcal{E})$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\leq\alpha$, which implies that for any $S$ not containing $i$, $H_{0,S}(\mathcal{E})$ must be rejected. So for any $R$ such that $H_{0,R}(\mathcal{E})$ is not rejected, we must have $i\in R$, which implies that $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$. ∎
Now we prove Proposition 2.
@@
If there exists some $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected, that is, $p_{R}>\alpha$. Then $\tilde{p}^{*}_{i}(\alpha)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\alpha}\}=\max\{p^{*}_{i},0\}=p^{*}_{i}$, so $\tilde{S}(\mathcal{E})=\widehat{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$ by Lemma A.1. Now we consider the remaining case that $H_{0,S}(\mathcal{E})$ is rejected for all $S\subseteq[m]$, that is, $p_{S}\leq\alpha$ for all $S\subseteq[m]$. Then, we have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ by definition, and $\tilde{p}^{*}_{i}(\alpha)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\alpha}\}=\max\{p^{*}_{i},1\}=1>\alpha$ for any $i\in[m]$. Hence $\tilde{S}(\mathcal{E})=\emptyset=\widehat{S}^{\text{ICP}}(\mathcal{E})$. ∎
A.3 Proof of Proposition 3
Proposition A.3.

### v3

@@ 4.2 Simultaneous true discovery bound for ICP
for each $S\subseteq[m]$. To this end, we generalize the p-value (5) and propose the following p-value for testing $H^{*}_{0,S}$:
@@
The validity of $p^{*}_{S}$ is shown in the following proposition.
@@
With valid p-values $p^{*}_{S}$ in hand, we can implement the closed testing procedure to obtain simultaneous true discovery bounds. A key observation here is that due to the special construction of $p^{*}_{S}$, there is no need to implement closed testing adjustments. To see this, if $p^{*}_{S}\leq\alpha$, we must have $p^{*}_{S’}\leq\alpha$ for all $S\subseteq S’$ ’, indicating that if a hypothesis is locally rejected, it must be rejected by closed testing. On the other hand, by definition, if a hypothesis is rejected by closed testing, it must be locally rejected. Therefore, a hypothesis is rejected by closed testing if and only if it is locally rejected, hence no closed testing adjustment is needed in this case. This leads to the following simultaneous true discovery bound:
$t_{\alpha}(R)=\min\{|R\setminus I|:I\subseteq R\text{ and }p^{*}_{I}>\alpha\}.$ (10)
In words, $t_{\alpha}(R)$ is the size of the smallest subset of $R$ whose corresponding complement in $R$ has a p-value larger than $\alpha$. It can be seen as a shortcut of the standard closed testing procedure, which simplifies the computation by not requiring the full implementation of all the steps in closed testing.
The theorem below formally shows that $t_{\alpha}(R)$ is a simultaneous true discovery bound.
@@
$P(|R\cap S^{*}|\geq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$ (11)
@@
Based on the simultaneous guarantee (11), users can freely check any set $R$ they regard as interesting, and calculate a high-probability lower bound $t_{\alpha}(R)$ on the true discoveries.
The simultaneous true discovery bound encompasses both FWER control and the defining sets (Heinze-Deml et al.,, 2018) as special cases. Specifically, by searching for the largest set $R$ whose $t_{\alpha}(R)=|R|$, we obtain a set with FWER control; by searching for all the smallest sets $R$ with $t_{\alpha}(R)=1$, we obtain the defining sets. In particular, it is easy to check that for $\widehat{S}^{\text{ICP}}$, we have $t_{\alpha}(\widehat{S}^{\text{ICP}})=|\widehat{S}^{\text{ICP}}|$, thereby recovering the FWER control guarantee of ICP. Hence, the same causal discoveries and information obtained from the original ICP approach are fully retained. This indicates that using our proposed simultaneous true discovery bound to check more sets is always beneficial, which will only extract additional useful information from data. See Section 5 for illustrations.
In fact, one can also obtain a simultaneous true discovery bound based on the original ICP formulation, without resorting to the multiple testing framework  [^1] We thank an anonymous referee for pointing out this.. Specifically, let
$\tilde{t}_{\alpha}(R)=\underset{S\subseteq[m]\text{ with }p_{S}>\alpha}{\min}|R\cap S|,\quad\text{for all }R\subseteq[p].$ (12)
Then, we have
$P(|R\cap S^{*}|\geq\tilde{t}_{\alpha}(R)\text{ for any }R\subseteq[m])\geq P(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected})\geq 1-\alpha,$
where the first inequality holds by the definition of $\tilde{t}_{\alpha}(R)$.
Although obtained based on different ideas, $t_{\alpha}(R)$ and $\tilde{t}_{\alpha}(R)$ (see (10) and (12)) are actually equivalent, as shown in the following theorem.
Theorem 4.2.
Let $\alpha\in(0,1)$. For any $R\subseteq[m]$,
$t_{\alpha}(R)=\tilde{t}_{\alpha}(R).$
It is not uncommon in the literature that two seemingly different methods are actually equivalent. For example, the simultaneous bound based on closed testing from Goeman and Solari, (2011) is equivalent to the bound from Genovese and Wasserman, (2006), which does not use closed testing (Goeman et al.,, 2021). Recognizing such connections is valuable because it provides different perspectives on the same method and allows insights to be gained from different angles. For example, in our case, understanding that $\tilde{t}_{\alpha}(R)$ is equivalent to a closed testing bound helps clarify its admissibility (Goeman et al.,, 2021), which is rather unclear from its original formulation (12).
@@ 5.1 Numerical simulations
We implement simulations to verify the simultaneous guarantee (11) and to illustrate the extra information one may obtain by using the simultaneous true discovery bound, compared to the original ICP approach. All simulations were carried out in R, and the code is available at https://github.com/Jinzhou-Li/ICPsimultaneousBounds.
@@
$X^{e}\leftarrow BX^{e}+\epsilon^{e}+b^{e},$
where $X^{e}=(X^{e}_{1},\dots,X^{e}_{10})^{T}$, $B$ is a strict lower-triangular matrix, $\epsilon^{e}\sim N_{10}(0,\text{diag}(\sigma))$ and $e\in\{1,2,3,4,5\}$. That is, we consider mean-shift intervention and five environments. We treat $X_{6}$ as the response in our simulations.
We implement $500$ simulations. In each simulation, we randomly sample the non-zero elements of matrix $B$ with a non-zero probability $0.8$. Their values are uniformly sampled from $(1,2)$. Each entry of $\sigma$ is uniformly sampled from $(0.5,1.5)$. For the mean-shift intervention, we set $b^{1}=0$. For other $b^{i}$, we randomly select $3$ non-zero entries excluding $X_{6}$. That is, we consider $3$ intervened variables. The non-zero entries of $b^{i}$ are sampled from $N(10,5^{2})$. We generate $100$ samples for each environment and use a significance level $\alpha=0.05$.
For each generated dataset, we apply ICP and record the number of discoveries. Based on the same p-values used by ICP, we calculate the simultaneous true discovery upper bound $t_{\alpha}(R)$ (see (10)) for all $5122^{9}=$ 512 sets. We also present the false discovery lower bounds $|R|-t_{\alpha}$ for illustration purposes, which essentially convey the same information as the simultaneous true discovery bound.
@@
Over these $500$ simulations, the empirical probability of $P(|R\cap S^{*}|\geq t_{\alpha}(R)\text{ for any }R\subseteq[m])$ is $0.988$, which is larger than $0.951-\alpha=$ 0.95, so the simultaneous guarantee (11) holds. The average number of causal predictors is $4.144$, and the average number of discoveries of ICP is $0.52$. The first plot in Figure 1 shows the average false discovery upper bounds and true discovery lower bounds for all $512$ sets, as well as the size of each set. Note that the x-axis is only the index for each set, and we didn’t present which sets these indices refer to for simplicity. Compared to ICP, extra information can be obtained by looking at these bounds. For example, by looking at the true discovery lower bound of the set containing all predictors (with set index $512$), we know that, on average, there are about three true causal predictors in all nine predictors.
To better illustrate the usage of the simultaneous bounds in practice, we look at two simulated datasets rather than the averaged results. These results are shown in the second and third plots of Figure 1, respectively.
For the first simulated dataset, the true causal predictors are $\{X_{1},X_{2},X_{3},X_{4}\}$. ICP discovers $X_{2}$ and $X_{3}$, meaning that with a probability larger than $0.95$, $X_{2}$ and $X_{3}$ are causal predictors. That is all the information we can get by using ICP. By looking at the simultaneous bounds (see the second plot in Figure 1), however, extra information can be obtained. For example, we can see that the sets with indices $131$ and $257$ has true discovery lower bounds $3$ and $4$, respectively. These two sets are $\{X_{1},X_{2},X_{3},X_{4}\}$ and $\{X_{1},X_{2},X_{3},X_{4},X_{5}\}$. Hence, we know that with probability larger than $0.95$, there are at least three causal predictors in $\{X_{1},X_{2},X_{3},X_{4}\}$, and at least four causal predictors in $\{X_{1},X_{2},X_{3},X_{4},X_{5}\}$. Note that the same information that both $X_{2}$ and $X_{3}$ are true causal predictors can also be obtained, because the true discovery lower bound for $\{X_{2},X_{3}\}$ (with set index $13$) is two.
@@ 5.2 Real data application
We look at a real dataset used in Stock et al., (2003) about the educational attainment of teenagers in the US (Rouse,, 1995). This dataset contains information of $4739$ students from approximately 1100 US high schools, including gender, ethnicity, achievement test score, whether the father or mother is a college graduate, whether the family owns their home, whether the school is in an urban area, county unemployment rate, state hourly wage in manufacturing, average college tuition, whether the family income above 25000 per year, and region.
@@
We first apply ICP with a significance level $\alpha=0.05$, returning one variable ‘score’ as the causal predictor. It is reasonable that test score has a causal influence on whether the student would attain a BA degree or higher. We then calculate the simultaneous bounds for all $81922^{13}=$ 8192 sets, and the results are shown in Figure 2. More information can be obtained by looking at these bounds. For example, by looking at the set with index $112$, we know that with a probability larger than $0.95$, there are at least two causal predictors in {‘score’, ’fcollege__\__no’, ’mcollege__\__no’}. It seems plausible that the true discovery lower bound is $2$ for this set, because the two variables of whether father or mother went to college are highly correlated, so it is difficult to distinguish between whether both are causal predictors or only one of them is.
6 Discussion
Motivated by the issue that FWER control in the original ICP approach can be overly conservative for some applications, resulting in few or no causal discoveries and the loss of valuable causal information, we address the error control problem of ICP in this paper. We focus on two commonly used and more liberal error control guarantees: FDR control and simultaneous true discovery bounds. We investigate whether it is possible to generalize the original ICP approach to possess these two guarantees, thereby revealing more causal information. The main idea is to reformulate ICP within the framework of multiple testing. We propose valid p-values for this multiple testing problem and show the equivalence of this formulation to the original ICP approach. Our answer for FDR control problem is negative: in this special multiple testing problem where the p-values are constructed based on ICP, FDR seems not a suitable error criterion. However, for simultaneous true discovery bounds, we provide a positive answer and propose such bounds based on closed testing. Through simulations and a real dataset, we demonstrate that using simultaneous true discovery bounds can indeed uncover more valuable causal information compared to the original ICP approach. Importantly, this additional information comes at no cost, as the original information is fully retained, making it always advantageous to use in practice.
The original ICP method has been generalized to various settings, including non-linear models (Heinze-Deml et al.,, 2018), sequential data (Pfister et al.,, 2019), and transformation models (Kook et al.,, 2023). It would be interesting to either directly apply or generalize our proposed simultaneous true discovery bound in these settings.
When testing the null hypothesis $H^{*}_{0,i}$ (see (4)), p-values based on the ICP approach are constructed by testing for invariance across different environments. However, other sources of information, such as testing for conditional independence, may also provide evidence to reject $H^{*}_{0,i}$. If new p-values can be obtained by combining information from different principles when testing $H^{*}_{0,i}$, then the simultaneous false discovery bound based on closed testing can be directly applied to provide causal discoveries with error control guarantees, and FDR control may become suitable again. Note that it is unclear how to achieve this with the original ICP formulation. This is another advantage of the multiple testing formulation (4). For future research, it would be interesting to develop concrete methods for combining information based on different principles when testing $H^{*}_{0,i}$ to achieve error-controlled causal discoveries.
Finally, we note that calculating the simultaneous true discovery bound (10) can become computationally intractable for large $|R|$ (or large $m$). This computational challenge is also inherent in the original ICP approach, which requires testing $2^{|m|}$ hypotheses. Various strategies have been proposed to address this issue, including reducing the set of potential causal predictors through pre-screening methods such as Lasso (see Peters et al., (2016) for further discussion). Addressing this computational challenge is another interesting direction for future research.
@@ Acknowledgement
Jinzhou Li thanks Nicolai Meinshausen for helpful discussions and suggestions on the draft of this paper. Jinzhou Li gratefully acknowledges support from the SNSF Grant P500PT-210978.
SUPPLEMENTARY MATERIAL
Appendix A A brief recap on closed testing
The closed testing procedure was originally introduced for controlling the familywise error rate (Marcus et al.,, 1976). Later, Goeman and Solari, (2011) proposed simultaneous false discovery bounds based on it. Below, we present the three main steps of the closed testing procedure.
Consider a multiple testing problem with hypotheses $H_{1},\dots,H_{m}$. For any $I\subseteq[m]$, let $H_{I}=\cap_{i\in I}H_{i}$ be an intersection hypothesis, where $H_{I}$ is true if and only if $H_{i}$ is true for all $i\in I$. The closed testing procedure proceeds as follows:
  • 1.
    Obtain local test results: Locally test all intersection hypotheses $H_{I}$ for $I\subseteq[m]$ at level $\alpha\in(0,1)$.
  • 2.
    Implement the closed testing adjustment: $H_{I}$ is rejected by closed testing if and only if $H_{J}$ is locally rejected in Step 1 for all $J\supseteq I$.
  • 3.
    Obtain simultaneous false discovery bound $u^{ct}(R)$: For any set $R\subseteq[m]$, let $u^{ct}(R)$ be the size of the largest subset of $R$ that is not rejected by closed testing in Step 2.
Then, Goeman and Solari, (2011) proved that $u^{ct}(R)$ is a simultaneous false discovery bound. That is,
${\mathbb{P}}(|R\cap\mathcal{N}|\leq u^{ct}(R),\forall R\subseteq[m])\geq 1-\alpha,$
where $\mathcal{N}\subseteq[m]$ is the set of null hypotheses.
Equivalently, $t^{ct}(R)=|R|-u^{ct}(R)$ is a simultaneous true discovery bound satisfying
${\mathbb{P}}(|R\cap S^{*}|\geq t^{ct}(R),\forall R\subseteq[m])\geq 1-\alpha,$
where $S^{*}=[m]\setminus\mathcal{N}$ is the set of true hypotheses. In particular, by the definition of $u^{ct}(R)$, we have that $t^{ct}(R)$ is the smallest subset of $R$ whose complement in $R$ is not rejected by closed testing.
Appendix B Proofs
B.1 Proof of Proposition 3.1
@@ Appendix B Proofs
(Proposition 3.1) When $H^{*}_{0,i}$ is true, we have $S^{*}\subseteq[m]\setminus\{i\}$ as $i\not\in S^{*}$, so $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{S^{*}}$. Therefore, for any $c\in(0,1)$,
@@
B.2 Proof of Proposition 3.2
To prove Proposition 3.2, we first introduce the following lemma, which gives the relation between $\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}\leq\alpha\}$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$.
Lemma B.1.
Assume that there exists $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected. Let $\widehat{S}(\mathcal{E})=\{i:p^{*}_{i}\leq\alpha\}$. Then, $\widehat{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$.
We first prove Lemma B.1.
@@
(Lemma B.1) If $\widehat{S}(\mathcal{E})=\emptyset$, that is, $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}>\alpha$ for all $i\in[m]$, then we must have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$. Otherwise, if there exists some $j\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, then for any $R$ whose corresponding $p_{R}>\alpha$, we must have $j\in R$ by the definition of $\widehat{S}^{\text{ICP}}(\mathcal{E})$. But $p^{*}_{j}=\max_{S\subseteq[m]\setminus\{j\}}p_{S}>\alpha$ implies that there exists some $R^{\prime}$ not containing $j$ and $p_{R^{\prime}}>\alpha$, which leads to a contradiction.
If $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$, by the assumption that there exists $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected (i.e., $p_{R}>\alpha$), there are two possibilities: (1) $R=\emptyset$ is the only set whose corresponding $p_{R}>\alpha$. In this case, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\geq p_{R}>\alpha$ for all $i\in[m]$, so $\widehat{S}(\mathcal{E})=\emptyset$. (2) There exist two different sets $R_{1}$ and $R_{2}$ such that $p_{R_{1}}>\alpha$, $p_{R_{2}}>\alpha$ and $R_{1}\cap R_{2}=\emptyset$. Then we must have $\widehat{S}(\mathcal{E})=\emptyset$. Otherwise, if there exists some $j\in\widehat{S}(\mathcal{E})$, that is, $p^{*}_{j}=\max_{S\subseteq[m]\setminus\{j\}}p_{S}\leq\alpha$, then in order for $p_{R_{1}}>\alpha$ and $p_{R_{2}}>\alpha$, both $R_{1}$ and $R_{2}$ must contain $j$, which contradicts the fact that $R_{1}\cap R_{2}=\emptyset$.
Hence, $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ if and only if $\widehat{S}(\mathcal{E})=\emptyset$.
Now assume both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ are non-empty.
If $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$, that is, for any $S$ such that $H_{0,S}(\mathcal{E})$ is not rejected, we have $i\in S$. So for any $R\subseteq[m]\setminus\{i\}$, $H_{0,R}(\mathcal{E})$ must be rejected, so $p^{*}_{i}=\max_{R\subseteq[m]\setminus\{i\}}p_{R}\leq\alpha$, which implies that $i\in\widehat{S}(\mathcal{E})$.
If $i\in\widehat{S}(\mathcal{E})$, we have $p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}\leq\alpha$, which implies that for any $S$ not containing $i$, $H_{0,S}(\mathcal{E})$ must be rejected. So for any $R$ such that $H_{0,R}(\mathcal{E})$ is not rejected, we must have $i\in R$. Combining with the assumption that there exists $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected, we have $i\in\widehat{S}^{\text{ICP}}(\mathcal{E})$.
Combining the above cases, we have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\widehat{S}(\mathcal{E})$. ∎
Now we prove Proposition 3.2.
@@
(Proposition 3.2) If there exists some $R\subseteq[m]$ such that $H_{0,R}(\mathcal{E})$ is not rejected, that is, $p_{R}>\alpha$. Then $\tilde{p}^{*}_{i}(\alpha)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\alpha}\}=\max\{p^{*}_{i},0\}=p^{*}_{i}$, so $\tilde{S}(\mathcal{E})=\widehat{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$, where the second equality is by Lemma B.1.
Now we consider the remaining case that $H_{0,S}(\mathcal{E})$ is rejected for all $S\subseteq[m]$, that is, $p_{S}\leq\alpha$ for all $S\subseteq[m]$. Then, we have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ by definition, and $\tilde{p}^{*}_{i}(\alpha)=\max\{p^{*}_{i},\Pi_{S\subseteq[m]}\mathbbm{1}_{p_{S}\leq\alpha}\}=\max\{p^{*}_{i},1\}=1>\alpha$ for any $i\in[m]$. Hence $\tilde{S}(\mathcal{E})=\emptyset=\widehat{S}^{\text{ICP}}(\mathcal{E})$. ∎
B.3 Proof of Proposition 3.3